# Fluid mechanics, Applied mathematics, Mathematical modelling

### v2

@@ 4 Linear stability
The model provides a powerful means to analyze band formation in ways that experiment and DNS do not have access to. In particular, with the model we can directly address the transition from uniform turbulence to turbulent bands via linear stability analysis. We develop this approach over the next two sections. In this section, we presents numerical results and establish the fidelity of long-wavelength approximation to the linear stability equations. In the next section, we derive a bound on the critical angle for the onset of patterns in the long-wavelength approximation.
@@ 4.1 Stability analysis of the full system
Recall that the uniform turbulent state in the model is a steady state of the form $(u_{0},u_{1},v_{1},w_{0},w_{1},q_{0})=(0,u_{ss},0,0,0,q_{ss})$, where $u_{ss}$ and $q_{ss}$ satisfy equations (20) after dropping the $t$ derivatives. Straightforward linearization of the model equations about a uniform turbulent state gives an eigenvalue problem for linear modes of the form $\hat{u}_{0}\exp(ik_{x}x+ik_{z}z+\sigma t)$ and similarly for $\hat{u}_{1},\hat{v}_{1},\hat{w}_{0},\hat{w}_{1},\hat{q}_{0}$, where $\sigma$ is the temporal growth rate and $k_{x}$ and $k_{z}$ are streamwise and spanwise wavenumbers of the perturbation. The linearized model equations are stated in full in appendix B and details of their derivation are found in the electronic supplementary material [55].
For a given value of $Re$ and pair of wavenumbers $(k_{x},k_{z})$, the linearized equations become a simple matrix eigenvalue problem that can be written
@@
Stability of the uniform states is indicated in the bifurcation diagram of figure 6b. The laminar state is linearly stable for all $Re$, independently of $(k_{x},k_{z})$. The lower-branch turbulent states are unstable. The focus is on possible instability of the upper-branch turbulent states. At large $Re$ the upper branch is linearly stable, as expected since uniform turbulence is the observed state in full model simulations With decreasing $Re$, a positive real eigenvalue first appears at a critical Reynolds number $Re_{c}=85.1$. Hence, for $Re<Re_{c}$ uniform turbulence is unstable to spatial modulations, as indicated by the dashed portion of the upper branch in figure 6b, and as seen in figures 7c and 8f.
@@
Figure 10 shows the maximum growth rates in the $(k_{x},k_{z})$ plane for two values of $Re$: one just above the saddle-node marking the appearance of the upper branch and one just below $Re_{c}$ where instability sets in. A mode with wavenumbers $(k_{x},k_{z})$ corresponds to a stripped pattern with wavelength $\lambda=2\pi/k=2\pi/\sqrt{k_{x}^{2}+k_{z}^{2}}$ and angle $\theta$ to the streamwise direction, where $\tan\theta=k_{x}/k_{z}$. At the onset of instability, $Re=Re_{c}$, the critical values of the wavelength and angle are: $\lambda_{c}=23.6$ and $\theta_{c}=22.5^{\circ}$. With decreasing $Re$, the unstable region enlarges and the fastest growing mode shifts to larger wavelength and slightly lower angle ($\lambda\approx 32$ and $\theta\approx 21.5^{\circ}$ at $Re=73$). These results align very well with evidence of a linear instability of uniform turbulence in plane Poiseuille flow at a critical angle of $\theta_{c}\approx 23^{\circ}$ [19]. Nonlinear model simulations of the 1D tilted domain setup show that the bifurcation to bands at $Re_{c}$ and $\theta_{c}$ is supercritical. That is, modulation strength grows continuously from zero and the transition is not hysteretic. The bifurcation is not always supercritical, however; for some angles and wavelengths the transition from uniform turbulence to bands is subcritical. This is will discussed in a separate publication.
@@ 4.2 Long-wavelength simplification
Formally, we assume $k_{x}/\beta\ll 1$ and $k_{z}/\beta\ll 1$, or for simplicity, $k/\beta\ll 1$, so that terms order $(k/\beta)^{2}$ terms are negligible compared with order $(k/\beta)$, and larger, terms. For example, the three-dimensional spectral Laplacian $-(k^{2}+\beta^{2})$ is approximated as $-\beta^{2}$. We relegate the details to appendix B, and to §5.2 where we take up the stability equations explicitly. At present, it is sufficient to understand that the long-wavelength approximation slightly alters the stability matrix $\bf M$ in equation (21). We do not introduced separate notation for the stability matrix, nor eigenvalues and vectors, for the approximate system.
Making the long-wavelength approximation, the critical values for the onset of patterns change slightly as shown in table 1. The dominant effect of the approximation is a shift in the critical Reynolds number to a slightly larger value. Figure 10c shows the leading eigenvalue of the approximate system for the same conditions as in figure 10b. The region of instability in $(k_{x},k_{z})$ is slightly larger under the approximation, which can be accounted for by the shift in $Re_{c}$, but otherwise plots 10b and 10c are very similar. Table 2 compares all six eigenvalues at typical parameter value of interest, demonstrating that the simplifying approximation preserves the full character of the linear stability problem. We now proceed to analyse the transition to patterns in the simplified stability equations.
@@ 5 Analysis of the critical point
We show explicitly that, in this long-wavelength approximation, the selected angle of turbulent bands at onset, the critical angle $\theta_{c}$, is bounded by $|\theta_{c}|<45^{\circ}$.
@@ 5.1 Infinitesimal balances
${\rm wall~{}normal:}$ 00 0 $=\underbrace{-(\alpha+\beta^{2}/{Re})\hat{v}_{1}}_{\rm dissipation}\underbrace{-ik_{x}A^{\prime}_{ss}\hat{q}_{0}}_{\rm Re.~{}stress},$ (22c)
${\rm TKE:}$ 00 0 $=\frac{1}{2}\beta A_{ss}\hat{u}_{1}+\left.\frac{\partial\dot{q}_{0}}{\partial q_{0}}\right|_{u_{ss},q_{ss}}\hat{q}_{0}.$ (22d)
@@
[…]
full $85.1$ $23.6$ $22.5^{\circ}$
long wavelength $93.3$ $20.1$ $21.5^{\circ}$
Table 1: Critical values for the full eigenvalue problem and simplified, long-wavelength, eigenvalue problem.
[…]
full $-0.05972$ $-0.03587\pm 0.04809i$ $-0.02906\pm 0.07571i$ 0.00069
long wavelength $-0.05968$ $-0.03505\pm 0.04809i$ $-0.02843\pm 0.07622i$ 0.00160
Table 2: Eigenvalues for the full eigenvalue problem and simplified, long-wavelength, eigenvalue problem for $Re=82$, $k_{x}=0.1$, $k_{z}=0.24$.
@@
Terms on the left-hand-side result from linearization of the advective nonlinearity about the homogeneous steady state $(u_{0},u_{1},v_{1},w_{0},w_{1},q_{0})=(0,u_{ss},0,0,0,q_{ss})$. Factors of $1/2$ in (22a) result from Galerkin projection onto the mean. The pressure term enforces incompressibility of the infinitesimal flow perturbation and is calculated in the electronic supplementary material [55]. Although the spanwise flow components do not appear in equations (22), their presence is necessary for determining the pressure term. The dissipation terms, viscous and Ekman friction, are straightforward as these are linear effects. The Reynold stresses are modelled with $A(q_{0})$, and when linearized about $q_{ss}$ become $A^{\prime}_{ss}\hat{q}_{0}$, where $A^{\prime}_{ss}\equiv dA/dq(q_{ss})$. The Reynold-stress forces in (22b) and (22c) include a pressure projection to divergence-free form.
@@
We have separated the $k$-independent terms from those proportional to $-k^{2}$, and used that $k_{x}^{2}=k^{2}\sin^{2}\theta$. The three terms in $L_{0}$ are the linearization of the production, the Ekman friction, and the dissipation, respectively, where $\varepsilon^{\prime}=d\varepsilon_{0}/dq_{0}(q_{ss})$. These are necessarily the terms that occur when differentiating (20b) with respect to $q_{0}$, and they are functions only of $Re$, and not wavenumbers.
@@ 5.2 Angle selection
From the marginal stability equations, we derive a bound on the band angle at the onset of patterns. All calculations follow exactly equations (22), and no further approximations are made. The algebra is straightforward, but we relegate full details to the electronic supplementary material [55].
@@
We explicitly indicate the dependencies on $k$, $\theta$, and $Re$.
Letting $F_{0}(k,\theta,Re)$ denote the right-hand side of (24), the condition for a zero eigenvalue, $\sigma=0$, is written $0=F_{0}(k,\theta,Re)$. This is simply the statement that the determinant of the linear stability matrix is zero at a bifurcation, but written with readily identifiable physical meanings for each term. We are interested in the critical condition where a zero eigenvalue first appears with decreasing $Re$. At such a point we much have
@@
where $Re_{c}$ is the critical Reynolds number and $(k_{c},\theta_{c}$) are the associated critical wavenumber and critical angle  [^1] To understand why (26c) must hold, suppose that $\partial_{\theta}F_{0}\neq 0$ at the critical point $(k_{c},\theta_{c},Re_{c})$. Then by the implicit function theorem, there is a function $\theta=g(Re)$ such that $F_{0}(k_{c},g(Re),Re)=0$ for $Re$ on an open interval containing $Re_{c}$. This implies that there is a zero eigenvalue both above and below $Re_{c}$. The contradicts the assumption that $Re_{c}$ is the maximum $Re$ for which there is a zero eigenvalue. A similar argument applies fixing $\theta$ and varying $k$.. In principle, given the model closure parameters, the three expressions (26) could be evaluated and used to find $Re_{c}$, $k_{c}$, and $\theta_{c}$. We have not found this to be analytically viable. However, the condition on the derivative with respect to angle (26c) itself can be used to establish a bound on $\theta_{c}$.
@@
One sees clearly that this expression arises because the advection and pressure terms in (24) have different dependencies on $\theta$. Taking into account the $\theta$-dependence of $L_{+}$, the exact statement is instead the inequality (see electronic supplementary material [55]):
@@ 6 Conclusion
First, we have obtained a model incorporating the large-scale flow associated with transitional turbulence in a planar flow geometry. Advection plays a dominant role in transitional turbulence and the large-scale flow is an essential ingredient to faithful modelling. In earlier work on pipe flow [10, 28, 27], a single scalar variable (the mean shear or, equivalently, the centreline velocity) was shown to be sufficient to capture the main effects of the large-scale mean flow. This was possible because, in pipes, flow is confined transverse to the streamwise direction. In contrast, planar cases are unconstrained in the spanwise direction and the large-scale flows are correspondingly richer. This has held back development of planar models. By considering stress-free boundaries and a sinusoidal driving force, we have succeeded in capturing the incompressible, vector character of the large-scale flow using a minimal expansion in five fields.
@@
Our third contribution is new theoretical insights into turbulent bands – the key building block of transitional turbulence in a planar setting. In the model, periodic turbulent bands (turbulent-laminar patterns) unambiguously bifurcate via a linear instability of uniform turbulence as $Re$ is decreased. While this linear instability is expected given the literature on turbulent-laminar patterns [4], especially recent work by Kashyap et al. [19], the model captures the instability in a simple linear system for momentum and energy balances. Analysing this linear system we establish a selection criterion for the pattern orientation at onset, namely, $0<|\theta_{c}|<45^{\circ}$, where the critical angle $\theta_{c}$ is the angle of the pattern to the streamwise direction when instability first occurs with decreasing $Re$. Oblique periodic bands in planar shear flows have been ubiquitously observed at angles $45^{\circ}$ or smaller, often in the range $24^{\circ}$ to $30^{\circ}$, and never at $0^{\circ}$ [4, 12, 22, 24, 25, 56, 19, 13, 57, 58, 62, 21]. Our analysis provides physical insight into what selects such angles – the relationship between advection and pressure in the streamwise momentum. Moreover, the bound suggests a reason why turbulent structures (puffs) in pipes do not arise as periodic patterns from uniform turbulence [64]. Confinement forces the wavevector to point in the streamwise direction, corresponding to $\theta=90^{\circ}$, which is outside the allowed range of angles for instability to periodic patterns.
@@
Finally, the model opens several new avenues of research. One could add new mechanisms or modify the model to describe other flow configurations, in particular pressure-driven pipes and channels. One of the more important avenues to pursue is a detailed understanding of the transition scenario using dynamical systems theory, similar to the successful analysis of models of pipe flow [27, 29, 65]. Closely related to this would be to exploit the physical mechanisms captured by the model to understand how energy balances adjust and break down with decreasing $Re$ [9, 20]. As stated in §5.2, it should be possible to improve the bound on the angle selection by taking into account details of the turbulent field. Finally, the effect of turbulent fluctuations can be included via a noise term to investigate rare events [66, 67, 68] and percolation transitions [11, 12, 13, 14] in the planar setting.
@@
The code used to run the model simulations described in the main text is v1.0 of ‘2D MWF’ [69]. It is preserved at https://doi.org/10.5281/zenodo.15261063, available via the Creative Commons Attribution 4.0 International (CC-BY) license. The code depends largely on the Dedalus package in Python [44], an open-source pseudo-spectral solver for partial differential equations. The electronic supplementary material [55], preserved in the Figshare repository https://doi.org/10.6084/m9.figshare.28903058, contains (i) details of the stability analysis, bound on the critical angle and description of the supplementary movies, (ii) the supplementary movies, (iii) the data used for the figures in this work, and (iv) the scripts used to generate the figures in this work.

### v3

@@ 4 Linear stability
The model provides a powerful means to analyze band formation in ways that experiment and DNS do not have access to. In particular, with the model we can directly address the transition from uniform turbulence to turbulent bands via linear stability analysis. We develop this approach over the next two sections. In this section, we presents numerical results and establish the fidelity of long-wavelength approximation to the linear stability equations. In the next section, we derive a bound on the critical angle for the onset of patterns in the long-wavelength approximation. In this section, we refer to the approximate system as the ‘long wavelength’ system, and the model without any approximations as the ‘full’ system.
@@ 4.1 Stability analysis of the full system
Recall that the uniform turbulent state in the model is a steady state of the form $(u_{0},u_{1},v_{1},w_{0},w_{1},q_{0})=(0,u_{ss},0,0,0,q_{ss})$, where $u_{ss}$ and $q_{ss}$ satisfy equations (24) after dropping the $t$ derivatives. Straightforward linearization of the model equations about a uniform turbulent state gives an eigenvalue problem for linear modes of the form $\hat{u}_{0}\exp(ik_{x}x+ik_{z}z+\sigma t)$ and similarly for $\hat{u}_{1},\hat{v}_{1},\hat{w}_{0},\hat{w}_{1},\hat{q}_{0}$, where $\sigma$ is the temporal growth rate and $k_{x}$ and $k_{z}$ are streamwise and spanwise wavenumbers of the perturbation. The linearized model equations are stated in full in appendix B and details of their derivation are found in the electronic supplementary material [SM].
For a given value of ReRe and pair of wavenumbers $(k_{x},k_{z})$, the linearized equations become a simple matrix eigenvalue problem that can be written
@@
Stability of the uniform states is indicated in the bifurcation diagram of figure 6b. The laminar state is linearly stable for all ReRe, independently of $(k_{x},k_{z})$. The lower-branch turbulent states are unstable. The focus is on possible instability of the upper-branch turbulent states. At large ReRe the upper branch is linearly stable, as expected since uniform turbulence is the observed state in model simulations. With decreasing ReRe, a positive real eigenvalue first appears at a critical Reynolds number $Re_{c}=85.1$. Hence, for $Re<Re_{c}$ uniform turbulence is unstable to spatial modulations, as indicated by the dashed portion of the upper branch in figure 6b, and as seen in figures 7c and 8f.
@@
Figure 10 shows the maximum growth rates in the $(k_{x},k_{z})$ plane for two values of ReRe: one just above the saddle-node marking the appearance of the upper branch and one just below $Re_{c}$ where instability sets in. A mode with wavenumbers $(k_{x},k_{z})$ corresponds to a stripped pattern with wavelength $\lambda=2\pi/k=2\pi/\sqrt{k_{x}^{2}+k_{z}^{2}}$ and angle $\theta$ to the streamwise direction, where $\tan\theta=k_{x}/k_{z}$. At the onset of instability, $Re=Re_{c}$, the critical values of the wavelength and angle are: $\lambda_{c}=23.6$ and $\theta_{c}=22.5^{\circ}$. With decreasing ReRe, the unstable region enlarges and the fastest growing mode shifts to larger wavelength and slightly lower angle ($\lambda\approx 32$ and $\theta\approx 21.5^{\circ}$ at $Re=73$). These results align very well with evidence of a linear instability of uniform turbulence in plane Poiseuille flow at a critical angle of $\theta_{c}\approx 23^{\circ}$ [kashyap2022linear]. Nonlinear model simulations of the 1D tilted domain setup show that the bifurcation to bands at $Re_{c}$ and $\theta_{c}$ is supercritical. That is, modulation strength grows continuously from zero as ReRe is decreased below $Re_{c}$ with $\theta$ fixed at $\theta_{c}$. The transition is not hysteretic. The bifurcation in the model is not always supercritical, however; for some fixed angles and wavelengths, the transition from uniform turbulence to bands as ReRe is decreased is subcritical. This is will discussed in a separate publication.
@@ 4.2 Long-wavelength simplification
Formally, we assume $k_{x}/\beta\ll 1$ and $k_{z}/\beta\ll 1$, or for simplicity, $k/\beta\ll 1$, so that terms order $(k/\beta)^{2}$ terms are negligible compared with order $(k/\beta)$, and larger, terms. For example, the three-dimensional spectral Laplacian $(k^{2}+\beta^{2})$ is approximated as $\beta^{2}$. We relegate the details to appendix B, and to §5.2 where we take up the stability equations explicitly. At present, it is sufficient to understand that the long-wavelength approximation slightly alters the stability matrix $\bf M$ in equation (25). We do not introduced separate notation for the stability matrix, nor eigenvalues and vectors, for the approximate system.
Making the long-wavelength approximation, the critical values for the onset of patterns change slightly to: $Re_{c}=93.3$, $\lambda_{c}=20.1$ and $\theta_{c}=21.5^{\circ}$. The dominant effect of the approximation is a shift in the critical Reynolds number to a slightly larger value. Figure 10c shows the leading eigenvalue of the approximate system for the same conditions as in figure 10b. The region of instability in $(k_{x},k_{z})$ is slightly larger under the approximation, which can be accounted for by the shift in $Re_{c}$, but otherwise plots 10b and 10c are very similar. We now proceed to analyse the transition to patterns in the simplified stability equations.
@@ 5 Analysis of the critical point
We show explicitly that, in this long-wavelength approximation, the selected angle of turbulent bands at onset, the critical angle $\theta_{c}$, is bounded by $0<|\theta_{c}|<45^{\circ}$.
@@ 5.1 Infinitesimal balances
${\rm wall~normal:}$ $0$ $=\underbrace{-(\alpha+\beta^{2}/{Re})\hat{v}_{1}}_{\rm dissipation}\underbrace{-ik_{x}A^{\prime}_{ss}\hat{q}_{0}}_{\rm Re.~stress},$ (26c)
${\rm TKE:}$ $0$ $=\frac{1}{2}\beta A_{ss}\hat{u}_{1}+\left.\frac{\partial\dot{q}_{0}}{\partial q_{0}}\right|_{u_{ss},q_{ss}}\hat{q}_{0}.$ (26d)
@@
Terms on the left-hand-side result from linearization of the advective nonlinearity about the homogeneous steady state $(u_{0},u_{1},v_{1},w_{0},w_{1},q_{0})=(0,u_{ss},0,0,0,q_{ss})$. Factors of $1/2$ in (26a) result from Galerkin projection onto the mean. The pressure term enforces incompressibility of the infinitesimal flow perturbation and is calculated in the electronic supplementary material [SM]. Although the spanwise flow components do not appear in equations (26), their presence is necessary for determining the pressure term. The dissipation terms, viscous and Ekman friction, are straightforward as these are linear effects. The Reynold stresses are modelled with $A(q_{0})$, and when linearized about $q_{ss}$ become $A^{\prime}_{ss}\hat{q}_{0}$, where $A^{\prime}_{ss}\equiv dA/dq(q_{ss})$. The Reynold-stress forces in (26b) and (26c) include a pressure projection to divergence-free form.
@@
We have separated the $k$-independent terms from those proportional to $k^{2}$, and used that $k_{x}^{2}=k^{2}\sin^{2}\theta$. The three terms in $L_{0}$ are the linearization of the production, the Ekman friction, and the dissipation, respectively, where $\varepsilon^{\prime}=d\varepsilon_{0}/dq_{0}(q_{ss})$. These are necessarily the terms that occur when differentiating (24b) with respect to $q_{0}$, and they are functions only of ReRe, and not wavenumbers.
@@ 5.2 Angle selection
From the marginal stability equations, we derive a bound on the band angle at the onset of patterns. All calculations follow exactly equations (26), and no further approximations are made. The algebra is straightforward, but we relegate full details to the electronic supplementary material [SM].
@@
We explicitly indicate the dependencies on $k$, $\theta$, and ReRe.
Letting $F_{0}(k,\theta,Re)$ denote the right-hand side of (28), the condition for a zero eigenvalue, $\sigma=0$, is written $0=F_{0}(k,\theta,Re)$. This is simply the statement that the determinant of the linear stability matrix is zero at a bifurcation, but written with readily identifiable physical meanings for each term. We are interested in the critical condition where a zero eigenvalue first appears with decreasing ReRe. At such a point we much have
@@
where $Re_{c}$ is the critical Reynolds number and $(k_{c},\theta_{c}$) are the associated critical wavenumber and critical angle  [^1] To understand why (30c) must hold, suppose that $\partial_{\theta}F_{0}\neq 0$ at the critical point $(k_{c},\theta_{c},Re_{c})$. Then by the implicit function theorem, there is a function $\theta=g(Re)$ such that $F_{0}(k_{c},g(Re),Re)=0$ for ReRe on an open interval containing $Re_{c}$. This implies that there is a zero eigenvalue both above and below $Re_{c}$. The contradicts the assumption that $Re_{c}$ is the maximum ReRe for which there is a zero eigenvalue. A similar argument applies fixing $\theta$ and varying $k$.. In principle, given the model closure parameters, the three expressions (30) could be evaluated and used to find $Re_{c}$, $k_{c}$, and $\theta_{c}$. We have not found this to be analytically viable. However, the condition on the derivative with respect to angle (30c) itself can be used to establish a bound on $\theta_{c}$.
@@
One sees clearly that this expression arises because the advection and pressure terms in (28) have different dependencies on $\theta$. Taking into account the $\theta$-dependence of $L_{+}$, the exact statement is instead the inequality (see electronic supplementary material [SM]):
@@ 6 Conclusion
First, we have obtained a model incorporating the large-scale flow associated with transitional turbulence in a planar flow geometry. Advection plays a dominant role in transitional turbulence and the large-scale flow is an essential ingredient to faithful modelling. In earlier work on pipe flow [barkley2011simplifying, barkley2015rise, barkley2016theoretical], a single scalar variable (the mean shear or, equivalently, the centreline velocity) was shown to be sufficient to capture the main effects of the large-scale mean flow. This was possible because, in pipes, flow is confined transverse to the streamwise direction. In contrast, planar cases are unconstrained in the spanwise direction and the large-scale flows are correspondingly richer. This has held back development of planar models. By considering stress-free boundaries and a sinusoidal driving force, we have succeeded in capturing the incompressible, vector character of the large-scale flow using a minimal expansion in five fields.
@@
Our third contribution is new theoretical insights into turbulent bands – the key building block of transitional turbulence in a planar setting. In the model, periodic turbulent bands (turbulent-laminar patterns) unambiguously bifurcate via a linear instability of uniform turbulence as ReRe is decreased. While this linear instability is expected given the literature on turbulent-laminar patterns [tuckerman2020patterns], especially recent work by Kashyap et al. [kashyap2022linear], the model captures the instability in a simple linear system for momentum and energy balances. Analysing this linear system we establish a selection criterion for the pattern orientation at onset, namely, $0<|\theta_{c}|<45^{\circ}$, where the critical angle $\theta_{c}$ is the angle of the pattern to the streamwise direction when instability first occurs with decreasing ReRe. Oblique periodic bands in planar shear flows have been ubiquitously observed at angles $45^{\circ}$ or smaller, often in the range $24^{\circ}$ to $30^{\circ}$, and never at $0^{\circ}$ [tuckerman2020patterns, chantry_universal, prigent2003long, duguet2010formation, duguet2013oblique, ShimizuPRF2019, kashyap2022linear, klotz2022phase, xiong2015turbulent, xiao2020growth, kashyap2020flow, reetz2019exact]. Our analysis provides physical insight into what selects such angles – the relationship between advection and pressure in the streamwise momentum. Moreover, the bound suggests a reason why turbulent structures (puffs) in pipes do not arise as periodic patterns from uniform turbulence [moxey2010distinct]. Confinement forces the wavevector to point in the streamwise direction, corresponding to $\theta=90^{\circ}$, which is outside the allowed range of angles for instability to periodic patterns.
@@
Finally, the model opens several new avenues of research. One could add new mechanisms or modify the model to describe other flow configurations, such as pressure-driven pipes and channels, and rotating flows [meseguer2009instability, berghout2020direct, feldmann2023routes, tsukahara2010flow]. One of the more important avenues to pursue is a detailed understanding of the transition scenario using dynamical systems theory, similar to the successful analysis of models of pipe flow [barkley2016theoretical, frishman2022dynamical, frishman2022mechanism]. Closely related to this would be to exploit the physical mechanisms captured by the model to understand how energy balances adjust and break down with decreasing ReRe [gome2023patterns1, gome2023patterns2]. As stated in §5.2, it should be possible to improve the bound on the angle selection by taking into account details of the turbulent field. Finally, the effect of turbulent fluctuations can be included via a noise term to investigate rare events [rolland2018extremely, rolland_2022, gome2022extreme] and percolation transitions [lemoult2016directed, chantry_universal, klotz2022phase, hof2023directed] in the planar setting.
@@
The code used to run the model simulations described in the main text is v1.0 of ‘2D MWF’ [benavides2025code]. It is preserved at https://doi.org/10.5281/zenodo.15261063, available via the Creative Commons Attribution 4.0 International ($C$-BY) license. The code depends largely on the Dedalus package in Python [burns2020dedalus], an open-source pseudo-spectral solver for partial differential equations. The electronic supplementary material [SM], preserved in the Figshare repository https://doi.org/10.6084/m9.figshare.28903058, contains (i) details of the stability analysis, bound on the critical angle and description of the supplementary movies, (ii) the supplementary movies, (iii) the data used for the figures in this work, and (iv) the scripts used to generate the figures in this work.